Translate the problem situation to a system of equations. Do not attempt to solve.
The perimeter of a rectangular building is 234 feet. The width is 25 feet shorter than the length. What are the dimensions? (Let W represent the width and L represent the length.)

Answer by richwmiller(17219)

Here we are just asked to set up the equations and not solve them.
The perimeter of a rectangular building is 234 feet. The width is 25 feet shorter than the length. What are the dimensions? (Let W represent the width and L represent the length.)
the formula for the perimeter of a rectangle is 2L+2W=Perimeter
we know that W=L-25
the perimeter =234
and that 2L+2W=234
So now we have two equations with two variables.
There is enough information to solve them.
We can substitute L-25 in the last equation for every W